The charge moving along a wire is governed by the charge equation q = 3t sin (30πt) C, determine the current at t = 300 μs
1
Expert's answer
2021-09-12T19:07:47-0400

"q=3tsin(30\pi t)\\i=\frac{dq}{dt}"

"i=\frac{d}{dt}(3tsin(30\pi t))"


"i=3sin(30\pi t)+90\pi t cos(30\pi t)"

Put t value


"i=3sin(30\pi \times300\times 10^{-6})+90\pi \times300\times 10^{-6}cos(30\pi\times300\times10^{-6} )"

"i=0.08481+0.08474=0.1695Amp"